two objects travelling around two different circular orbits with constant speed they both have the same acceleration but object a is travelling twice as object b.the orbit radius for object a is? and the orbit radius for object b


1
Expert's answer
2020-12-04T07:40:23-0500

Solution

Relation between acceleration and velocity and radius

"a=\frac{v^2}{r}"

For object a

"a_a=\frac{v_a^2}{r_a}"

For object b

"a_b=\frac{v_b^2}{r_b}"

Given relation

"v_a=2v_b"

Therefore using given condition

"a_a=a_b"

"\frac{4v_b^2}{r_a}=\frac{v_b^2}{r_b}"

"r_a=4r_b"

This is relation between both radius.